OpenStudy (kirbykirby):

Given the following pmf (prob. mass function), find its cdf: \[f(x)=\frac{7}{10}(\frac{3}{10})^x, x=0,1,2,...\]

OpenStudy (kirbykirby):

I know the formula to find the cdf is \[F(x)=\sum_{X_j \le X}^{}f(x_j)\] Does this mean you get \[\sum_{j=0}^{\infty}\frac{7}{10}(\frac{3}{10})^{X_j}\]? I dunno why I'm slightly confused about the discrete case. I was used to doing this for continuous rv's.